# LiDMaS: Architecture-Level Modeling of Fault-Tolerant Magic-State Injection in GKP Photonic Qubits

### v1

@@ Abstract
Fault-tolerant quantum computation in photonic architectures relies critically on the efficient preparation of high-fidelity logical magic states under realistic constraints on finite squeezing and photon loss. In this work, we investigate logical $T$-gate magic-state preparation in GKP-encoded photonic qubits using a repeat-until-success (RUS) injection protocol combined with outer surface-code protection. We adopt an architecture-level modeling framework based on a custom lightweight density-matrix simulator implemented using standard numerical linear algebra, in which finite squeezing is mapped to effective logical dephasing, depolarizing noise is included at the logical level, and photon loss is treated as a heralded erasure process. This approach deliberately avoids explicit continuous-variable wavefunction simulation, hardware-specific photonic models, and quantum software frameworks such as PennyLane, enabling transparent and computationally efficient exploration of architectural trade-offs. We systematically explore squeezing values from $8$ to $16$ dB, baseline loss probabilities in the range 0.010.010.030.03, and surface-code distances $d=1,3,5,$ and $7$. Across this parameter space, we evaluate RUS success probability, average injection overhead, and logical magic-state fidelity. We find that the RUS protocol achieves success probabilities exceeding $0.94$ across all studied regimes, with an average overhead of approximately 1.151.151.201.20 rounds per successful injection. After outer-code protection, logical magic-state fidelities reach $F_{\mathrm{log}}\approx 0.77$$0.80$, exhibiting weak sensitivity to moderate photon loss but a strong monotonic dependence on squeezing. Sensitivity analysis confirms that finite squeezing is the dominant continuous error source, while loss primarily affects heralded failure rates rather than logical quality. Phase-boundary diagrams further identify the minimum squeezing required to simultaneously satisfy success-probability thresholds ($\geq 0.95$) and logical-fidelity targets ($\geq 0.79$) as a function of code distance. These results provide quantitative design guidelines for scalable, fault-tolerant photonic quantum architectures.
@@ I Introduction
At the architectural level, surface-code-based protection has been proposed as a powerful outer layer to suppress residual logical errors in photonic systems. Notably, Nickerson, Li, and Benjamin [15] demonstrated that heralded photon loss can be efficiently tolerated by erasure-aware decoding, significantly relaxing loss requirements for large-scale quantum computation. However, their analysis focused primarily on Clifford operations and logical error rates, leaving open the question of how non-Clifford gate synthesis, and in particular magic-state fidelity, behaves under realistic loss and noise conditions [7]. Understanding this behavior is essential for assessing the viability of photonic fault-tolerant architectures beyond threshold arguments.
@@
In this work, we address these open questions by presenting a systematic, architecture-level study of logical $T$-gate magic-state preparation in GKP-encoded photonic qubits. Rather than performing explicit continuous-variable wavefunction simulations or decoder-level modeling, we adopt a deliberately simplified yet physically informed framework in which logical qubits are represented by $2\times 2$ density matrices and noise processes are modeled as effective logical channels. Finite GKP squeezing is mapped to an effective Pauli-$Z$ dephasing channel, residual imperfections are captured by logical depolarizing noise, and photon loss is treated as a heralded erasure process that aborts failed attempts without degrading successful outcomes. This abstraction allows us to efficiently explore a wide parameter space while retaining direct correspondence with experimentally relevant quantities.
We focus on a repeat-until-success (RUS) magic-state injection protocol combined with an outer surface-code abstraction, enabling us to quantify key operational metrics, including success probability within a finite round cap, average injection overhead, and logical magic-state fidelity after outer-code protection. By sweeping squeezing levels, baseline loss probabilities, and surface-code distances, we construct sensitivity maps and phase-boundary diagrams that reveal how hardware quality and error correction jointly determine performance. Importantly, this approach allows us to disentangle the roles of squeezing and loss: while loss primarily affects heralded failure rates and overhead, finite squeezing emerges as the dominant factor limiting logical fidelity.
The main contributions of this work are threefold. First, we provide a unified architecture-level model that explicitly incorporates finite squeezing, depolarizing noise, and heralded photon loss into logical magic-state preparation, without relying on CV simulation frameworks or quantum software libraries such as PennyLane. Second, we demonstrate through sensitivity analysis that logical magic-state fidelity is remarkably robust to moderate loss once heralded erasure is accounted for, but remains strongly dependent on squeezing across all studied code distances. Third, we introduce quantitative phase-boundary diagrams that identify the minimum squeezing required to simultaneously achieve high success probability and high logical fidelity as a function of loss and code distance, offering practical design guidance for photonic quantum computing platforms.
@@ II.1 Architecture-Level Modeling Framework
The results presented in this work are obtained using a custom, lightweight density-matrix simulation framework designed to capture the dominant logical error mechanisms relevant to fault-tolerant photonic quantum architectures [2, 6]. Rather than performing explicit continuous-variable wavefunction simulations of Gottesman–Kitaev–Preskill (GKP) states or relying on quantum software frameworks such as PennyLane [1], we adopt an architecture-level approach in which encoded qubits are represented by $2\times 2$ density matrices and logical operations are modeled using standard matrix algebra. This deliberate abstraction enables efficient numerical exploration of design trade-offs while preserving the essential structure of logical noise propagation and error correction [16].
Finite squeezing of approximate GKP states is incorporated through an effective logical dephasing channel, parameterized by a squeezing-dependent error probability calibrated to reproduce the expected monotonic improvement with increasing squeezing [8, 11]. Additional logical depolarizing noise is included to model residual non-Pauli errors arising from imperfect Clifford operations and finite stabilizer-cycle fidelity [10]. Photon loss, the dominant hardware impairment in photonic platforms, is treated as a heralded erasure process that terminates the current magic-state injection attempt without contributing unheralded logical errors [15]. This modeling choice reflects the strong loss-detectability inherent to photonic architectures and allows the separation of failure probability from logical-state quality.
@@ II.2 Logical noise model
We model these effects using the composite map in Equation 1
@@ II.2.1 Finite-squeezing-induced dephasing
Finite-energy GKP states exhibit residual phase uncertainty that manifests as effective logical dephasing after GKP error correction. We model this as a Pauli-$Z$ channel (Equation 2),
@@
To relate this to the experimental squeezing parameter $s$ (expressed in dB), we introduce a monotonic proxy mapping in Equation 3
$p_{Z}(s)=\min!\mathopen{}\mathclose{{\left(0.5,,\alpha_{s}e^{-\beta s}}}\right),$ (3)
@@ II.3 Logical $T$-gate magic-state injection
We consider the standard magic-state injection protocol based on the $|A\rangle$ state in Equation 6,
@@ II.3.1 Injection circuit
We designate one measurement branch as the successful $T$-gate realization up to Clifford equivalence, while the complementary branch is treated as unsuccessful and triggers repetition. This choice reflects a standard architecture-level abstraction used in RUS analyses.
@@ II.4 Outer-code abstraction: surface-code protection
Successful injection produces a logical state that is further protected by an outer surface code of distance $d$. Rather than explicitly simulating stabilizer cycles and decoding, we adopt a standard scaling-law abstraction for the logical error rate (Equation 9):
@@
where $p_{\mathrm{th}}$ is the threshold error rate and $A$ is a prefactor of order unity.
@@
where $w_{Z}$ and $w_{P}$ weight residual dephasing and depolarizing contributions after GKP correction and stabilizer cycles.
@@ II.5 Sensitivity analysis
To quantify robustness, we compute finite-difference gradients of the logical fidelity with respect to loss and squeezing (Equation 12):
@@ II.6 Phase-boundary construction
Finally, we identify threshold-like operating regions by imposing joint constraints in Equation 13;
@@
For each baseline loss and code distance, we compute the minimum squeezing $s_{\min}$ satisfying both conditions, yielding a phase boundary in Equation 14;
@@ II.7 Numerical implementation
All simulations were performed using Monte Carlo sampling of density-matrix evolution under the effective noise channels described above. Each data point corresponds to $5\times 10^{3}$ independent RUS trials, ensuring statistical convergence of success probabilities and fidelities. Random seeds were fixed for reproducibility.
@@ III.1 Architecture-level performance of GKP-based magic-state injection
We begin by evaluating the performance of a repeat-until-success (RUS) logical $T$-gate injection protocol implemented using GKP-encoded photonic qubits and protected by an outer surface code. The analysis is conducted at the architecture level, incorporating effective noise models for finite squeezing, photon loss, and logical depolarization, while avoiding explicit decoder or syndrome simulation.
Figure 3 shows the RUS success probability as a function of the squeezing proxy for several baseline loss values and surface-code distances. Across all loss regimes considered, the success probability increases monotonically with squeezing and remains high ($\gtrsim 0.9$) even at moderate squeezing levels. Increasing the surface-code distance has only a weak effect on the success probability itself, indicating that the RUS dynamics are primarily governed by the injection circuit and heralded loss rather than by outer-code suppression.
Figure 3: RUS success probability as a function of the squeezing proxy for different baseline loss values and surface-code distances. Higher squeezing consistently improves the probability of successful magic-state injection, while the dependence on code distance remains weak.
@@ III.2 RUS overhead and repeat-until-success behavior
To quantify the operational overhead of the injection protocol, we next examine the average number of RUS rounds required to achieve a successful injection, conditioned on success. The results are shown in Fig. 4.
The average number of rounds remains close to unity across the entire parameter space, typically between $1.15$ and $1.20$. Notably, the overhead exhibits only a weak dependence on squeezing and loss, reflecting the fact that unsuccessful branches are efficiently heralded and restarted. Increasing the surface-code distance does not significantly increase the RUS overhead, demonstrating that logical protection does not introduce prohibitive latency in the injection process.
Figure 4: Average number of RUS rounds required for successful magic-state injection as a function of squeezing. The overhead remains close to unity across loss regimes and code distances, indicating efficient heralding and fast convergence of the RUS protocol.
@@ III.3 Logical magic-state fidelity after outer-code protection
For all loss values considered, the logical fidelity improves monotonically with squeezing, reflecting the reduction of effective GKP-induced phase noise. Increasing the surface-code distance systematically enhances the logical fidelity, with diminishing returns at higher distances, consistent with the expected scaling of surface-code logical error rates. Importantly, once outer-code protection is applied, the logical fidelity becomes largely insensitive to moderate variations in loss, indicating that photon loss is effectively converted into heralded erasure rather than uncorrectable logical errors.
Figure 5: Average logical magic-state fidelity conditioned on successful injection, as a function of squeezing. Increasing the surface-code distance systematically improves logical fidelity, while moderate loss variations have a weak effect once error correction is applied.
@@ III.4 Sensitivity analysis
To quantify the robustness of logical magic-state preparation and to identify dominant error sources, we performed a sensitivity analysis of the logical fidelity with respect to both baseline photon loss and squeezing. Finite-difference gradients of the logical fidelity were evaluated over the discrete architecture-level parameter grid explored in the simulations.
Figures 6 present sensitivity heatmaps for surface-code distances $d=1,3,5,$ and $7$. Across all distances, the sensitivity of logical fidelity with respect to baseline loss remains close to zero throughout most of the parameter space, indicating strong robustness against moderate variations in photon loss. This behavior reflects the heralded nature of loss in photonic architectures, which primarily impacts success probability rather than the quality of successful logical states.
@@
Figure 6: Sensitivity heatmaps of logical magic-state fidelity with respect to baseline photon loss (left column: panels (a), (c), (e), (g)) and squeezing (right column: panels (b), (d), (f), (h)) for surface-code distances $d=1,3,5,$ and 77, arranged from top to bottom. Across all code distances, the sensitivity to loss remains near zero throughout the parameter space, reflecting the heralded nature of photon loss in photonic architectures and its limited impact on successful logical outcomes. In contrast, sensitivity to squeezing is pronounced at low squeezing values, indicating finite-energy GKP noise as the dominant continuous error mechanism. Increasing squeezing and code distance both reduce sensitivity, demonstrating enhanced robustness under outer-code protection.

### v3

@@ Abstract
Fault-tolerant quantum computation in photonic architectures requires efficient preparation of high-fidelity logical magic states under realistic constraints of finite squeezing and photon loss. We present LiDMaS (Lightweight Density-Matrix Simulator), an architecture-level study of logical $T$-gate magic-state preparation in Gottesman–Kitaev–Preskill (GKP)-encoded photonic qubits using a repeat-until-success (RUS) injection protocol combined with outer surface-code protection. A density-matrix simulator based on standard numerical linear algebra is employed, mapping finite squeezing to effective logical dephasing, incorporating logical depolarizing noise, and treating photon loss as a heralded erasure process. Parameter sweeps are performed over squeezing values from $8$ to $16$ dB, baseline loss probabilities between $0.005$ and $0.03$, and surface-code distances $d=1,3,5,$ and $7$. Across this regime, RUS success probabilities range from about $0.90$ to $0.99$, with average injection overhead between $1.15$ and $1.21$ rounds per successful attempt. After outer-code protection, logical fidelities reach $F_{\mathrm{log}}\approx 0.765\text{--}0.796$, exhibiting weak sensitivity to moderate photon loss but strong monotonic dependence on squeezing. Sensitivity analysis identifies finite squeezing as the dominant continuous error source, while loss primarily impacts heralded failure rates. Phase-boundary diagrams determine minimum squeezing requirements to achieve success probability $\geq 0.95$ and logical fidelity $\geq 0.79$ as a function of code distance, providing quantitative design guidance for scalable photonic fault-tolerant architectures.
@@ I Introduction
At the architectural level, surface-code-based protection has been proposed as a powerful outer layer to suppress residual logical errors in photonic systems. Notably, Nickerson, Fitzsimons, and Benjamin [15] demonstrated that heralded photon loss can be efficiently tolerated by erasure-aware decoding, significantly relaxing loss requirements for large-scale quantum computation. However, their analysis focused primarily on Clifford operations and logical error rates, leaving open the question of how non-Clifford gate synthesis, and in particular magic-state fidelity, behaves under realistic loss and noise conditions [7]. Understanding this behavior is essential for assessing the viability of photonic fault-tolerant architectures beyond threshold arguments.
@@
In this work, these open questions are addressed via a systematic architecture-level investigation of logical $T$-gate magic-state preparation in GKP-encoded photonic qubits. We refer to this framework as LiDMaS (Lightweight Density-Matrix Simulator). Rather than performing explicit continuous-variable wavefunction simulations or decoder-level modeling, we adopt a deliberately simplified yet physically informed framework in which logical qubits are represented by $2\times 2$ density matrices and noise processes are modeled as effective logical channels. Finite GKP squeezing is mapped to an effective Pauli-$Z$ dephasing channel, residual imperfections are captured by logical depolarizing noise, and photon loss is treated as a heralded erasure process that aborts failed attempts without degrading successful outcomes. This abstraction allows us to efficiently explore a wide parameter space while retaining direct correspondence with experimentally relevant quantities. The novelty is the explicit coupling of a RUS non-Clifford injection workflow to a photonic noise model with heralded loss and an outer-code abstraction, enabling direct estimates of success probability, overhead, and logical fidelity rather than only asymptotic thresholds.
Focus is placed on a repeat-until-success (RUS) magic-state injection protocol combined with an outer surface-code abstraction, enabling quantification of key operational metrics, including success probability within a finite round cap, average injection overhead, and logical magic-state fidelity after outer-code protection. By sweeping squeezing levels, baseline loss probabilities, and surface-code distances, sensitivity maps and phase-boundary diagrams are constructed to reveal how hardware quality and error correction jointly determine performance. Importantly, this approach disentangles the roles of squeezing and loss: while loss primarily affects heralded failure rates and overhead, finite squeezing emerges as the dominant factor limiting logical fidelity.
The main contributions of this work are threefold. First, a unified architecture-level model is provided that explicitly incorporates finite squeezing, depolarizing noise, and heralded photon loss into logical magic-state preparation, without relying on continuous-variable simulation frameworks or quantum software libraries such as PennyLane. Second, sensitivity analysis demonstrates that logical magic-state fidelity is remarkably robust to moderate loss once heralded erasure is accounted for, but remains strongly dependent on squeezing across all studied code distances. Third, quantitative phase-boundary diagrams are introduced to identify the minimum squeezing required to simultaneously achieve high success probability and high logical fidelity as a function of loss and code distance, thereby offering practical design guidance for photonic quantum computing platforms.
@@ II.1 Architecture-Level Modeling Framework
The results presented in this work are obtained using a custom, lightweight density-matrix simulation framework designed to capture the dominant logical error mechanisms relevant to fault-tolerant photonic quantum architectures [2, 6]. Rather than performing explicit continuous-variable wavefunction simulations of Gottesman–Kitaev–Preskill (GKP) states or relying on quantum software frameworks such as PennyLane [1], an architecture-level approach is adopted in which encoded qubits are represented by $2\times 2$ density matrices and logical operations are modeled using standard matrix algebra.This deliberate abstraction enables efficient numerical exploration of design trade-offs while preserving the essential structure of logical noise propagation and error correction [16].
Finite squeezing of approximate GKP states is incorporated through an effective logical dephasing channel, parameterized by a squeezing-dependent error probability calibrated to reproduce the expected monotonic improvement with increasing squeezing [8, 11]. Additional logical depolarizing noise is included to model residual non-Pauli errors arising from imperfect Clifford operations and finite stabilizer-cycle fidelity [10]. Photon loss, the dominant hardware impairment in photonic platforms, is treated as a heralded erasure process that terminates the current magic-state injection attempt without contributing unheralded logical errors [15]. This modeling choice reflects the strong loss-detectability inherent to photonic architectures and allows the separation of failure probability from logical-state quality. In implementation, all noise probabilities are clipped to the physical interval $[0,1]$, and the squeezing-to-dephasing proxy is further capped at $0.5$ to preserve complete positivity. The resulting logical noise model is summarized in Fig. 1.
@@ II.2 Logical noise model
These effects are modeled using the composite map in Equation 1.
@@ II.2.1 Finite-squeezing-induced dephasing
Finite-energy GKP states exhibit residual phase uncertainty that manifests as effective logical dephasing after GKP error correction. This effect is modeled as a Pauli-Z channel (Equation 2),
@@
To relate this to the experimental squeezing parameter $s$ (expressed in dB), a monotonic proxy mapping is introduced in Equation 3.
$p_{Z}(s)=\min\!\mathopen{}\mathclose{{\left(0.5,\,\alpha_{s}e^{-\beta s}}}\right),$ (3)
@@ II.3 Logical $T$-gate magic-state injection
The standard magic-state injection protocol based on the $|A\rangle$ state in Equation 6 is considered.
@@ II.3.1 Injection circuit
One measurement branch is designated as the successful $T$-gate realization up to Clifford equivalence, while the complementary branch is treated as unsuccessful and triggers repetition. This choice reflects a standard architecture-level abstraction used in RUS analyses.
@@ II.3.2 Repeat-until-success abstraction
Input: Squeezing values $\{s\}$, loss bases $\{p_{\mathrm{base}}\}$, distances $\{d\}$, trials NN, round cap $R_{\max}$
Output: Metrics: $P_{\mathrm{succ}}$, $\langle R\rangle_{\mathrm{succ}}$, $F_{\mathrm{inj}}$, $F_{\mathrm{log}}$
foreach $p_{\mathrm{base}}\in\{p_{\mathrm{base}}\}$ do
 foreach $s\in\{s\}$ do
    $p_{Z}\leftarrow\min(0.5,\,\alpha_{s}e^{-\beta s})$;
    $p_{E}^{\mathrm{data}}\leftarrow\mathrm{clip}\!\mathopen{}\mathclose{{\left(p_{\mathrm{base}}(1+\alpha_{\mathrm{LS}}p_{Z})}}\right)$;
    $p_{E}^{\mathrm{anc}}\leftarrow\mathrm{clip}\!\mathopen{}\mathclose{{\left(1.5\,p_{\mathrm{base}}(1+\alpha_{\mathrm{LS}}p_{Z})}}\right)$;
    foreach $d\in\{d\}$ do
         Initialize counters and accumulators;
       for $i=1$ to NN do
          for $r=1$ to $R_{\max}$ do
               Apply logical noise to data and ancilla; if erasure then abort trial;
               Apply CNOT, measure ancilla in $X$ basis, apply Clifford feedforward;
               Apply post-noise; if erasure then abort trial;
             if success branch then
                  Compute $F_{\mathrm{inj}}$ against $T|\psi\rangle$;
                $p_{\mathrm{phys}}\leftarrow w_{Z}p_{Z}+w_{P}p_{\mathrm{dep}}$;
                $p_{L}\leftarrow A\mathopen{}\mathclose{{\left(\frac{p_{\mathrm{phys}}}{p_{\mathrm{th}}}}}\right)^{(d+1)/2}$, capped at $1/2$ if $p_{\mathrm{phys}}\geq p_{\mathrm{th}}$;
                $F_{\mathrm{log}}\leftarrow(1-p_{L})F_{\mathrm{inj}}+p_{L}/2$;
                  Record success and break;
        Compute $P_{\mathrm{succ}}$, $\langle R\rangle_{\mathrm{succ}}$, and averages;
Algorithm 1 Architecture-level RUS $T$-state injection sweep
@@ II.4 Outer-code abstraction: surface-code protection
Successful injection produces a logical state that is further protected by an outer surface code of distance $d$. Rather than explicitly simulating stabilizer cycles and decoding, a standard scaling-law abstraction for the logical error rate is adopted (Equation 9):
@@
where $p_{\mathrm{th}}$ is the threshold error rate and $A$ is a prefactor of order unity. In implementation, if $p_{\mathrm{phys}}\geq p_{\mathrm{th}}$ the logical failure probability is pessimistically capped at $p_{L}=1/2$, corresponding to a maximally mixed logical output.
@@
where $w_{Z}$ and $w_{P}$ weight residual dephasing and depolarizing contributions after GKP correction and stabilizer cycles. In the numerical implementation, $w_{Z}=0.3$ and $w_{P}=0.1$, and $p_{\mathrm{phys}}$ is computed from the data-qubit noise channel only (ancilla and post-injection noise are not included in $p_{\mathrm{phys}}$).
@@ II.5 Sensitivity analysis
To quantify robustness, finite-difference gradients of the logical fidelity with respect to loss and squeezing are computed (Equation 12):
@@ II.6 Phase-boundary construction
Finally, threshold-like operating regions are identified by imposing joint constraints in Equation 13;
@@
For each baseline loss and code distance, the minimum squeezing $s_{\min}$ satisfying both conditions is computed, yielding a phase boundary in Equation 14;
@@ II.7 Numerical implementation
All simulations were performed using Monte Carlo sampling of density-matrix evolution under the effective noise channels described above. Each data point corresponds to $5\times 10^{3}$ independent RUS trials, ensuring statistical convergence of success probabilities and fidelities. Random seeds were fixed for reproducibility. A single fixed input state $|\psi\rangle$ (drawn once at initialization and normalized) is used across the full sweep to ensure comparability between parameter settings.
@@ III.1 Architecture-level performance of GKP-based magic-state injection
Evaluation begins with the performance of a repeat-until-success (RUS) logical T-gate injection protocol implemented using GKP-encoded photonic qubits and protected by an outer surface code.The analysis is conducted at the architecture level, incorporating effective noise models for finite squeezing, photon loss, and logical depolarization, while avoiding explicit decoder or syndrome simulation.
Figure 3 shows the RUS success probability as a function of the squeezing proxy for several baseline loss values and surface-code distances. Across all loss regimes considered, the success probability remains high ($\gtrsim 0.9$) and shows a weak upward trend with squeezing (within Monte Carlo fluctuations). Increasing the surface-code distance has only a weak effect on the success probability itself, indicating that the RUS dynamics are primarily governed by the injection circuit and heralded loss rather than by outer-code suppression.
Figure 3: RUS success probability as a function of squeezing for baseline loss values $p_{\mathrm{base}}=0.005$$0.03$ (panel titles). Each panel shows surface-code distances $d=1,3,5,7$; the success probability is high across the sweep and only weakly dependent on $d$. The vertical axis is zoomed to highlight small variations.
@@ III.2 RUS overhead and repeat-until-success behavior
To quantify the operational overhead of the injection protocol, the average number of RUS rounds required to achieve a successful injection, conditioned on success, is examined. The results are shown in Fig. 4.
The average number of rounds remains close to unity across the entire parameter space, typically between $1.15$ and $1.21$. Notably, the overhead exhibits only a weak dependence on squeezing and loss, reflecting the fact that unsuccessful branches are efficiently heralded and restarted. Increasing the surface-code distance does not significantly increase the RUS overhead, demonstrating that logical protection does not introduce prohibitive latency in the injection process.
Figure 4: Average number of RUS rounds required for successful magic-state injection, conditioned on success, versus squeezing. Each panel corresponds to a baseline loss value, and curves indicate surface-code distance. The overhead remains near unity throughout, and the vertical axis is zoomed to reveal small differences.
@@ III.3 Logical magic-state fidelity after outer-code protection
For all loss values considered, the logical fidelity generally improves with squeezing, reflecting the reduction of effective GKP-induced phase noise. Increasing the surface-code distance systematically enhances the logical fidelity, with diminishing returns at higher distances, consistent with the expected scaling of surface-code logical error rates. Importantly, once outer-code protection is applied, the logical fidelity becomes largely insensitive to moderate variations in loss, indicating that photon loss is effectively converted into heralded erasure rather than uncorrectable logical errors.
Figure 5: Average logical magic-state fidelity conditioned on successful injection as a function of squeezing. Panels sweep baseline loss and curves show surface-code distance; higher squeezing and larger $d$ increase logical fidelity, while moderate loss variations have a weak effect after outer-code protection. The vertical axis is zoomed to emphasize trends.
@@ III.4 Sensitivity analysis
To quantify the robustness of logical magic-state preparation and to identify dominant error sources, a sensitivity analysis of the logical fidelity with respect to both baseline photon loss and squeezing was performed. Finite-difference gradients of the logical fidelity were evaluated over the discrete architecture-level parameter grid explored in the simulations.
Figures 6 and 7 present sensitivity heatmaps for surface-code distances $d=1,3,5,$ and $7$. Across all distances, the sensitivity of logical fidelity with respect to baseline loss remains close to zero throughout most of the parameter space, indicating strong robustness against moderate variations in photon loss. This behavior reflects the heralded nature of loss in photonic architectures, which primarily impacts success probability rather than the quality of successful logical states.
@@
Figure 6: Sensitivity of logical fidelity to baseline photon loss, $\partial F_{\mathrm{log}}/\partial p_{\mathrm{base}}$, for surface-code distances $d=1,3,5,7$. Values remain near zero across the parameter space, indicating that heralded loss primarily affects success probability rather than the quality of successful logical states.
Figure 7: Sensitivity of logical fidelity to squeezing, $\partial F_{\mathrm{log}}/\partial s$, for surface-code distances $d=1,3,5,7$. Sensitivity is largest at low squeezing and diminishes with increasing squeezing and code distance, confirming finite-energy GKP noise as the dominant continuous error mechanism.